# Variable coherence model for free-electron laser pulses

### v1

@@
Abstract
@@
I Introduction
In recent years, free electron lasers (FELs) have enabled the generation of light pulses with intensities previously inaccessible through tabletop sources alone [1, 2, 3]. In particular, self-amplified spontaneous emission (SASE) [4] FELs can generate high-intensity sub-femtosecond soft x-ray pulses, thus enabling attosecond-science experiments targeting core-level electrons in atoms and molecules [5]. Recent FEL developments relevant for ultrafast science include pulse-pair generation that enable attosecond x-ray pump attosecond x-ray probe campaigns [1, 6], including two-color setups where each pulse has a different central photon energy [7, 8]. Ongoing efforts also aim at pulse shaping [9, 10] in the form of chirped x-rays pulses.
@@ I Introduction
In this paper, we revisit the partial coherence model of Ref. [24] to simulate SASE pulses with a controllable coherence width. Our variable coherence model (VCM) enables the generation of pulses that range from minimally coherent, like those within the partial coherence model, to maximally coherent akin to Fourier-limited pulses. The coherence width dictates the magnitude of the coherence across a pulse’s bandwidth, thereby allowing us to control its overall randomness while keeping the other parameters (central frequency and bandwidth) constant. Figure 1 illustrates the effect of introducing a variable coherence width, where we compare the spectral intensity profiles of three VCM pulses that only differ by their coherence widths. Clearly, increasing the coherence width results in pulses with fewer sub-pulses, leading to a single Gaussian shape when the coherence width is very large. The VCM also provides an avenue for including pulse shaping in the model while retaining controllable stochasticity within and between pulses.
Figure 1: Examples of VCM pulses of equation (1) generated using three different coherence widths $W$, as indicated in the title of each panel. Each pulse is generated using the FLASH parameter set in table 1.
We quantify the VCM properties with systematic statistical analyses and comparisons of the intensity and number of sub-pulses contained within pulses of varying coherence width. We also investigate the correlation between sub-pulse statistics in the time and frequency domains. We apply our VCM statistical analyses to three distinct FEL parameter sets loosely matching the LCLS-I, LCLS-II, and FLASH capabilities [24, 25, 26, 27] – see table 1 for the parameter values. Figure 2 shows an example of minimally coherent VCM pulses for each parameter set. Finally, we illustrate the effect of a varying coherence width on the reshaping of an absorption feature in a model atomic system.
Figure 2: Examples of typical VCM pulses with zero coherence width for each parameter set in table 1: long soft X-ray (LCLS-I), short soft X-ray (LCLS-II), and XUV (FLASH). In each panel, the red markers mark the sub-pulse peaks selected by our sub-pulse detection method – see section II.2.
@@
The structure of this paper is as follows: Section II introduces the methods we use for our analyses, including the VCM in subsection II.1 and the algorithm we use for sub-pulse detection in II.2. In section III we discuss the results of our statistical analyses, with sub-pulse analyses in the time and frequency domains in subsection III.1 as well as cross-correlation analyses of the time and frequency statistics in III.2. Section IV illustrates the effect of a varying coherence width on the reshaping of an absorption feature in a model atomic system. Finally, we conclude our findings in section V.
II Methods
In this section, we introduce the variable coherence model (VCM) and discuss its implementation for numerical simulations. Additionally, we establish the algorithms and criterion used for our sub-pulse analyses discussed in section III.
II.1 Variable coherence model
@@
where $\phi_{c}(\omega)=-\omega t_{0}$ is the Fourier-limited coherent piece and $\phi_{v}$ is randomly generated and responsible for the characteristic noise of SASE pulses. As pulse shaping capabilities like chirping become available in SASE FELs [28, 9, 10], one may incorporate such features in the coherent part of the phase $\phi_{c}$.
When the frequencies $\phi_{v}$ in equation (2) are sampled from a uniform random distribution over $0,2\pi]$ we recover the partial coherence model [24], which creates a minimally coherent pulse. Throughout this paper, we refer to this as the zero coherence width limit of the VCM. Figure 1 (a) shows an example of a zero coherence width pulse for the “FLASH” set of parameters of table 1. Alternatively, a non-zero coherence width restricts the phase variation across frequencies by sampling $\phi_{v}$ from a Lévy process. The coherence width then quantifies the rate of change in the phases, with small versus large coherence widths corresponding to fast versus slow changes between neighboring frequencies, respectively. We have experimented with two Lévy processes, a random walk and Brownian motion, and found very similar results between them. All the results we show in this paper are obtained with the Brownian motion case. In the limit of infinite coherence width, $\phi_{v}$ becomes a constant across all frequencies, and the resulting pulse is fully coherent. Figure 1 (b) shows an example of a VCM pulse with a nonzero coherence width. Compared to the zero coherence width case in panel (a), we see a clear reduction in the number of sub-pulses, reflecting the increased coherence between the pulse’s frequency components. In the limit of very large coherence width, as in panel (c), we observe a single broadband pulse with an envelope that closely matches the shape of the frequency window $W_{F}$ in equation (1). Small differences between $W_{F}$ and the observed pulse envelope come from the temporal window $W_{T}$ we impose on the overall pulse.
@@
II.2 Sub-pulse detection
While VCM pulses have a well-defined and smooth structure on average, the random phases $\phi_{v}$ in equation (2) yield individual pulses comprised of a random number of local peaks with drastically varying intensities; we refer to these peaks as sub-pulses. Figure 2 shows three examples of typical zero coherence width pulses generated across the three regimes of interest from table 1. For each pulse, the peaks of the constituent sub-pulses are identified with a red marker. In the time domain, VCM pulses are comprised of fast oscillations at the central frequency $\omega_{0}$ modulated by a slowly-varying envelope. In our sub-pulse analyses, we ignore these fast oscillations and only consider the slowly-varying envelopes. From figure 2, we identify that the number and intensity of sub-pulses are extremely sensitive to the pulse parameters used to create them. In section III, we investigate the statistical properties of the sub-pulses produced by the VCM, specifically regarding their dependence on the coherence width both in the time and frequency domains for each of the parameter sets in table 1.
To study the statistics associated with the sub-pulses, we first require a means for automatically capturing their locations and intensities within individual VCM pulses. The peak of each sub-pulse can be identified as a local maximum within the full pulse, which we obtain using the findpeaks function in MATLAB’s signal processing toolbox [29]. We note that the temporal and spectral windows $W_{T}$ and $W_{F}$ in equation (1) have the effect of drastically suppressing the intensities of sub-pulses that lie near the tails of any given pulse, so we discard these weaker sub-pulses from our statistical analyses. Specifically, we impose thresholds on the minimum peak heights and prominences. We use the prominence to distinguish the most relevant peaks based on how much they stand out compared to neighboring peaks. For all parameter sets listed in table 1 and both in the time and frequency domains, we disregard any sub-pulses that are less than 10% of the intensity of the most intense sub-pulse occurring within the same VCM pulse. Then, for all parameter sets in the frequency domain, we demand that the minimum peak prominence is at least 4.5% of the maximum sub-pulse intensity, whereas in the time domain we find that 10% is sufficient. Examples of qualifying sub-pulses are highlighted with red markers in figure 1. We have checked that our results are robust with small changes in the threshold criteria we use in our analyses.
@@ II.2 Sub-pulse detection
III Statistical Analyses
III.1 Sub-pulse statistics
@@
To begin our analyses, we examine the distributions of sub-pulse intensities in the frequency and time domains and show the results for the FLASH parameters in figure 3. Panel (a) shows the probability of generating a sub-pulse with some particular intensity as a function of the coherence width and the inset depicts a scaled view across a smaller range of values. For a low coherence width, below $\approx$ 0.25 eV or about half the pulses’ bandwidth, we see a large and non-uniform spread of possible sub-pulse intensities highlighted by the significant difference between the mean (dotted curve in the inset) and mode (solid) of the sub-pulse distribution. As the coherence width further increases, we see a narrowing of the distribution that converges towards the fully-coherent Fourier-limited case. This behavior at large coherence widths is precisely what we expect from a fully coherent pulse: one prominent peak with a consistent maximum intensity. In panel (b), we show the corresponding sub-pulse intensity distribution in the time domain. In this case, most of the sub-pulses at low coherence widths below $\approx$ 0.25 eV are concentrated towards smaller intensities rather than having a large spread as they did in the frequency domain – here the mean and mode (dotted and solid curves) of the sub-pulse distribution are close to each other. Then, as the coherence width is increased, we see that the most probable sub-pulse intensity gradually increases towards an asymptotic value just like it did in the frequency domain. Appendix A presents similar sub-pulse intensity analyses for the LCLS-I and LCLS-II sets of parameters. For LCLS-I, we see very similar behavior to the FLASH case, but the LCLS-II results are more distinct. Nonetheless, in all cases we observe a convergence towards asymptotic sub-pulse intensities.
@@
Figure 3: Sub-pulse intensity probability distributions as a function of the coherence width for the FLASH parameters of table 1 in the (a) frequency and (b) time domains. In each panel, the inset zooms in on the lower coherence widths where the probability transitions from low to high average value distributions. The dotted and solid curves mark the mean and mode of each distribution at a fixed coherence width, respectively. The vertical red line marks the bandwidth of the FLASH parameter set (0.55 eV).
@@
A notable difference between the time and frequency sub-pulse intensity distributions of figure 3 is how fast they converge towards the fully-coherent limit when the coherence width is increased. For a coherence width equal to the FEL bandwidth, we see in panel (a) that the distribution in the frequency domain is predominantly converged around its final value, whereas in the time domain, the distribution is still biased towards lower intensities until the coherence width is $\approx$ 1.5 eV, and not fully converged until the coherence width is near ten times the FEL bandwidth. Looking at individual VCM pulses, we attribute the difference between the time and frequency to how the pulses’ coherence width affects their structures: in the frequency domain, increasing the coherence width results in the sub-pulses progressively moving closer to the peak of $W_{F}$ from equation 1, naturally increasing their intensities – see for instance figure 1. In the time domain on the other hand, increasing the coherence width leads to a narrowing of the pulse around $t_{0}$, which means that most of the sub-pulses manifest as lower-intensity pedestals surrounding the narrow central peak. For intermediate coherence widths, these pedestals are very prevalent, and thus weigh down the most probable sub-pulse intensity, resulting in a slower transition towards the asymptotic intensity compared to the frequency domain.
@@
We investigate the effects of the coherence width on the other FEL parameter sets of table 1 in figure 4, where we show the average sub-pulse intensities and modes as a function of the coherence width for both the (a) frequency and (b) time domains – also shown as the dotted and solid curves in the insets of figure 3, respectively. For all curves, the lower and upper bounds of the shaded regions represent the 10th and $90^{th}$ percentiles of the distributions. In the frequency domain of panel (a), we recognize a similar behavior between the mean curves describing the FLASH and LCLS-I parameter sets while the LCLS-II case produces distinct results. For FLASH and LCLS-I, the average sub-pulse intensity first decreases when increasing the coherence width from zero. We attribute this feature to the large spread of possible intensities that initially weights the average towards a larger value at zero coherence width before clustering near lower intensities for small nonzero coherence widths. After reaching a shallow minimum, the average sub-pulse intensity converges towards its asymptotic Fourier-limited value, due to the concentration of sub-pulses towards the pulses’ central frequency. On the other hand, the LCLS-II average sub-pulse intensity is maximized at zero coherence width, followed by a slow but consistent decline until it reaches its asymptotic value. In the time domain of panel (b), we see that all parameter sets display similar qualitative behavior, with each average sub-pulse intensity curve monotonically increasing towards its asymptotic Fourier-limited value. Comparing the sub-pulse intensity distributions between the two panels (shaded areas), we see that broader distributions in the frequency domain are associated with narrower ones in the time domain, and vice versa, as a result of the uncertainty principle. We also see that the range of coherence widths required to reach the Fourier-limited asymptotic value is considerably larger in the time domain than in the frequency domain, as we observed in figure 3 for the FLASH case.
@@
Figure 4: Mean sub-pulse intensities (dotted) and sub-pulse modes (solid) as functions of the relative coherence width, defined as the ratio between the coherence width and average pulse bandwidth, for the LCLS-I (blue), LCLS-II (orange), and FLASH (purple) FEL parameter sets of table 1 in the (a) frequency and (b) time domains. For each data set, the upper and lower bounds of the shaded regions represent the 90th and 10th percentiles of the sub-pulse intensity distributions, respectively. In panel (a), we scaled the FLASH results for visual clarity.
@@
Next, we investigate the coherence width’s effects on the number of sub-pulses contained within a given VCM pulse. Figure 5 compares how the average number of sub-pulses depends on the coherence width across all parameter sets and in both the (a) frequency and (b) time domains. In all cases, we see that the average number of sub-pulses is maximized at zero coherence width. Then, we find that increasing the coherence width results in the average number of sub-pulses monotonically decreasing towards one, just like we would expect from a fully-coherent pulse. In the FLASH and LCLS-I cases, we further find that the decrease follows an exponential law. Like for the sub-pulse intensities of figure 4, we see that the distributions of sub-pulse numbers narrow with increasing coherence width – see the shaded areas at $\pm 1$ standard deviation around the average.
Figure 5: (a-b) Average number of sub-pulses (curves) plotted against the relative coherence width for the LCLS-I (blue), LCLS-II (orange), and FLASH (purple) FEL parameter sets of table 1 in the (a) frequency and (b) time domains. For each curve, the shaded regions represent $\pm$1 standard deviation around mean value. (c-f) Histograms of the sub-pulse number distributions for three exemplary values of the coherence width for the (c,d) LCLS-I (blue) and (e,f) FLASH (purple) parameter sets in the (c,e) frequency and (d,f) time domains.
@@
Figure 5 (c-f) shows individual sub-pulse number distributions at three exemplary values of the coherence width for the LCLS-I (top row) and FLASH (bottom row) parameter sets within both the frequency (left column) and time (right column) domains. At zero coherence width, we find that both parameter sets yield similar number distributions between the time and frequency domains, with a mean of $\approx 11$ sub-pulses for LCLS-I and $\approx 4$ sub-pulses for FLASH. As the coherence width increases well beyond the bandwidth of each parameter set, the distributions in the time domain are quicker to settle around unity. For each panel (c-f), the leftmost distributions correspond to the minimal coherence width required such that at least 90% of VCM pulses contain a single sub-pulse, thereby reliably producing pulses that are effectively coherent. For the LCLS-I parameter set, this coherence width is 28.8 eV ($\approx 5.2$ times the average pulse bandwidth) in the frequency domain and 23 eV ($\approx 4.2$ times) in the time domain, whereas for FLASH it is 3.57 eV ($\approx 6.5$ times) in the frequency domain and 2.61 eV ($\approx 4.7$ times) in the time domain. Similar to the sub-pulse intensity, we attribute the difference in the convergence rate for the sub-pulse number in time vs frequency to the effect of increasing the coherence with on the shapes of the pulses in the time domain (narrow central peak with pedestals) and frequency (limited effect on the average pulse envelope).
III.2 Joint Probability Distributions

### v2

@@
Variable coherence model for free-electron laser pulses
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
@@
1 Introduction
In recent years, free electron lasers (FELs) have enabled the generation of light pulses with intensities previously inaccessible through tabletop sources alone [1, 2, 3]. In particular, self-amplified spontaneous emission (SASE) [4] FELs can generate high-intensity sub-femtosecond soft x-ray pulses, thus enabling attosecond-science experiments targeting core-level electrons in atoms and molecules [5]. Recent FEL developments relevant for ultrafast science include pulse-pair generation that enable attosecond x-ray pump attosecond x-ray probe campaigns [1, 6], including two-color setups where each pulse has a different central photon energy [7, 8]. Ongoing efforts also aim at pulse shaping [9, 10] in the form of chirped x-ray pulses.
@@ 1 Introduction
In this paper, we revisit the partial coherence model of Ref. [24] to simulate SASE pulses with a controllable coherence width $\Omega$. Our variable coherence model (VCM) enables the simulation of pulses that range from minimally coherent, like those within the partial coherence model, to maximally coherent akin to Fourier-limited pulses. The coherence width dictates the magnitude of the coherence across a pulse’s bandwidth, thereby allowing us to control its overall randomness while keeping the other parameters (central frequency and bandwidth) constant. Figure 1 (a-c) illustrates the effect of introducing a variable coherence width, where we compare the spectral intensity profiles of three VCM pulses that only differ by their coherence widths. Clearly, increasing the coherence width results in pulses with fewer sub-pulses, leading to a single Gaussian shape when the coherence width is very large. The VCM also provides an avenue for including pulse shaping in the model while retaining controllable stochasticity within and between pulses.
Figure 1: Examples of VCM pulses of equation (1) simulated using (a-c) three different coherence widths $\Omega$ and the FLASH parameters set, and (d-f) zero coherence widths for each parameter set in table 1.
We quantify the VCM properties with systematic statistical analyses and comparisons of the intensity and number of sub-pulses contained within pulses of varying coherence width. We also investigate the correlation between sub-pulse statistics in the time and frequency domains. We apply our VCM statistical analyses to three distinct FEL parameter sets loosely matching the LCLS-I, LCLS-II, and FLASH capabilities [24, 25, 26, 27] – see table 1 for the parameter values. Figure 1 (d-f) shows an example of minimally coherent VCM pulses simulated for each parameter set. Finally, we illustrate the effect of a varying coherence width on the reshaping of an absorption feature in a model atomic system.
@@
The structure of this paper is as follows: Section 2 introduces the methods we use for our analyses, including the VCM in subsection 2.1 and the algorithm we use for sub-pulse detection in 2.2. In section 3 we discuss the results of our statistical analyses, with sub-pulse analyses in the time and frequency domains in subsection 3.1 as well as cross-correlation analyses of the time and frequency statistics in 3.2. Section 4 illustrates the effect of a varying coherence width on the reshaping of an absorption feature in a model atomic system. Finally, we conclude our findings in section 5.
2 Methods
In this section, we introduce the variable coherence model (VCM) and discuss its implementation for numerical simulations. Additionally, we establish the algorithms and criterion used for our sub-pulse analyses discussed in section 3.
2.1 Variable coherence model
@@
where $\phi_{c}(\omega)=-\omega t_{0}$ is coherent part of the phase, ensuring that the pulse converges towards the Fourier limit at infinite coherence width ($\Omega\to\infty$). $\phi_{v}$ contains the stochastic part of the pulse and is responsible for the characteristic noise of SASE pulses. As pulse shaping capabilities like chirping become available in SASE FELs [28, 9, 10], one may incorporate such features in the coherent part of the phase $\phi_{c}$.
When the phases $\phi_{v}$ in equation (2) are sampled from a uniform random distribution over $0,2\pi]$ we recover the partial coherence model [24], which creates a minimally coherent pulse. Throughout this paper, we refer to this as the zero coherence width limit ($\Omega=0$) of the VCM. Figure 1 (a,f) show examples of zero coherence width pulses for the “FLASH” set of parameters in table 1. We note that $\Omega=0$ still imposes some degree of coherence in the phases of neighboring frequencies, due to the convolution of the independent randomly selected $\phi(\omega)$ with the Fourier transform of the time window: $E(\omega)\propto\mathcal{F}[W_{T}]*(W_{F}e^{i\phi})$. Alternatively, a non-zero coherence width expands this local phase coherence across selectively larger portions of the spectrum. To do so, the VCM restricts the phase variation across frequencies by sampling $\phi_{v}$ from a Lévy process. Specifically, for a given frequency sampling $\omega_{1},\omega_{1},\ldots\omega_{n}$ we pick consecutive phases as
$\phi_{v}(\omega_{k+1})=\phi_{v}(\omega_{k})+\pi\sqrt{\frac{\omega_{k+1}-\omega_{k}}{\Omega}}\Phi,$ (3)
where $\Phi$ is a random variable with zero mean value and unity standard deviation. The scaling factor $\sqrt{\omega_{k+1}-\omega_{k}}$ ensures that the stochastic spread of the phases between different VCM pulses is independent of the frequency sampling. For a random walk $\Phi\sim\sqrt{12}\ \mathcal{U}(-0.5,0.5)$, with $\mathcal{U}$ the uniform random distribution, and for Brownian motion $\Phi\sim\mathcal{N}(0,1)$, with $\mathcal{N}$ the normal distribution. Finally, to ensure the randomization of the overall pulses’ carrier-envelope phase, we also add a global random phase sampled from $\mathcal{U}(0,2\pi)$ to all the $\phi_{v}$.
The coherence width $\Omega$ in equation (3) quantifies the stochasticity of the phase contribution from $\phi_{v}$, with smaller $\Omega$ (larger stochasticity) leading to more randomness within and between VCM pulses. We have experimented with both the random walk and Brownian motion and found very similar results between them. All the results we show in this paper are obtained with the Brownian motion case. In the limit of infinite coherence width, $\phi_{v}$ becomes a constant across all frequencies, and the resulting pulse is fully coherent. Figure 1 (b) shows an example of a VCM pulse with a nonzero coherence width. Compared to the zero coherence width case in panel (a), we see a clear reduction in the number of sub-pulses, reflecting the increased coherence between the pulse’s frequency components. In the limit of very large coherence width, as in panel (c), we observe a single broadband pulse with an envelope that closely matches the shape of the frequency window $W_{F}$ in equation (1). Small differences between $W_{F}$ and the observed pulse envelope come from the temporal window $W_{T}$ we impose on the overall pulse.
@@
2.2 Sub-pulse detection
While VCM pulses have a well-defined and smooth structure on average, the random phases $\phi_{v}$ in equation (2) yield individual pulses comprised of a random number of local peaks with drastically varying intensities; we refer to these peaks as sub-pulses. Figure 1 (d-f) shows three examples of typical zero coherence width pulses generated across the three regimes of interest from table 1. For each pulse, the peaks of the constituent sub-pulses are identified with a red marker. In the time domain, VCM pulses are comprised of fast oscillations at the central frequency $\omega_{0}$ modulated by a slowly-varying envelope. In our sub-pulse analyses, we ignore these fast oscillations and only consider the slowly-varying envelopes. From figure 1, we identify that the number and intensity of sub-pulses are extremely sensitive to the pulse parameters used to simulate them. In section 3, we investigate the statistical properties of the sub-pulses produced by the VCM, specifically regarding their dependence on the coherence width both in the time and frequency domains for each of the parameter sets in table 1.
To study the statistics associated with the sub-pulses, we first require a means for automatically capturing their locations and intensities within individual VCM pulses. The peak of each sub-pulse can be identified as a local maximum within the full pulse, which we obtain using the findpeaks function in MATLAB’s signal processing toolbox [29]. We note that the temporal and spectral windows $W_{T}$ and $W_{F}$ in equation (1) have the effect of drastically suppressing the intensities of sub-pulses that lie near the tails of any given pulse, so we discard these weaker sub-pulses from our statistical analyses by setting a threshold on the minimum peak heights. The stochastic nature of the pulse generation means that some sub-pulses may feature small local maxima within their temporal/spectral amplitude without these maxima being associated with a different sub-pulse. We similarly discard these local maxima by thresholding their prominence compared to neighboring peaks. Specifically, for all parameter sets listed in table 1 and both in the time and frequency domains, we disregard any sub-pulses that are less than 10% of the intensity of the most intense sub-pulse occurring within the same VCM pulse. Then, for all parameter sets, we demand that the peak prominence is at least 4.5% of the maximum sub-pulse intensity in the frequency domain, and 10% in the time domain. The larger prominence threshold in the time domain is chosen due to an increased amount of small local maxima within the sub-pulses’ envelopes. Examples of qualifying sub-pulses are highlighted with red markers in figure 1. We have checked that our results are robust with small changes in the threshold criteria we use in our analyses.
@@ 2.2 Sub-pulse detection
3 Statistical Analyses
3.1 Sub-pulse statistics
@@
To begin our analyses, we examine the distributions of simulated sub-pulse intensities in the frequency and time domains and show the results for the FLASH parameters in figure 2. Panel (a) shows the probability of generating a sub-pulse with some particular intensity as a function of the coherence width and the inset depicts a scaled view across a smaller range of values. For a low coherence width, below $\approx$ 0.25 eV or about half the pulses’ bandwidth, we see a large and non-uniform spread of possible sub-pulse intensities highlighted by the significant difference between the mean (dotted curve in the inset) and mode (solid) of the sub-pulse distribution. As the coherence width further increases, we see a narrowing of the distribution that converges towards the fully-coherent Fourier-limited case. This behavior at large coherence widths is precisely what we expect from a fully coherent pulse: one prominent peak with a consistent maximum intensity. In panel (b), we show the corresponding sub-pulse intensity distribution in the time domain. In this case, most of the sub-pulses at low coherence widths below $\approx$ 0.25 eV are concentrated towards smaller intensities rather than having a large spread as they did in the frequency domain – here the mean and mode (dotted and solid curves) of the sub-pulse distribution are close to each other. Then, as the coherence width is increased, we see that the most probable sub-pulse intensity gradually increases towards an asymptotic value just like it did in the frequency domain. The supplemental document presents similar sub-pulse intensity analyses simulated for the LCLS-I and LCLS-II sets of parameters. For LCLS-I, we see very similar behavior to the FLASH case, but the LCLS-II results are more distinct. Nonetheless, in all cases we observe a convergence towards asymptotic sub-pulse intensities.
@@
Figure 2: Simulated sub-pulse intensity probability distributions as a function of the coherence width for the FLASH parameters of table 1 in the (a) frequency and (b) time domains. In each panel, the inset zooms in on the lower coherence widths where the probability transitions from low to high average value distributions. The dotted and solid curves mark the mean and mode of each distribution at a fixed coherence width, respectively. The vertical red line marks the bandwidth of the FLASH parameter set (0.55 eV).
@@
A notable difference between the time and frequency sub-pulse intensity distributions of figure 2 is how fast they converge towards the fully-coherent limit when the coherence width is increased. For a coherence width equal to the FEL bandwidth, we see in panel (a) that the distribution in the frequency domain is predominantly converged around its final value, whereas for the time domain in (b), the distribution is still biased towards lower intensities until the coherence width is $\approx$ 1.5 eV, and not fully converged until the coherence width is near ten times the FEL bandwidth. Looking at individual VCM pulses, we attribute the difference between the time and frequency to how the pulses’ coherence width affects their structures: in the frequency domain, increasing the coherence width results in the sub-pulses progressively moving closer to the peak of $W_{F}$ from equation 1, naturally increasing their intensities – see for instance figure 1 (a-c). In the time domain on the other hand, increasing the coherence width leads to a narrowing of the pulse around $t_{0}$, which means that most of the sub-pulses manifest as lower-intensity pedestals surrounding the narrow central peak. For intermediate coherence widths, these pedestals are very prevalent, and thus weigh down the most probable sub-pulse intensity, resulting in a slower transition towards the asymptotic intensity compared to the frequency domain.
@@
We investigate the effects of the coherence width on the sub-pulse intensities simulated with the other FEL parameter sets of table 1 in figure 3 (a,b), where we show the average sub-pulse intensities and modes as a function of the coherence width for both the (a) frequency and (b) time domains – also shown as the dotted and solid curves in the insets of figure 2, respectively. For all curves, the lower and upper bounds of the shaded regions represent the 10th and $90^{th}$ percentiles of the distributions. In the frequency domain of panel (a), we recognize a similar behavior between the mean curves describing the FLASH and LCLS-I parameter sets while the LCLS-II case produces distinct results. For FLASH and LCLS-I, the average sub-pulse intensity first decreases when increasing the coherence width from zero. We attribute this feature to the large spread of possible intensities that initially weights the average towards a larger value at zero coherence width before clustering near lower intensities for small nonzero coherence widths. After reaching a shallow minimum, the average sub-pulse intensity converges towards its asymptotic Fourier-limited value, due to the concentration of sub-pulses towards the pulses’ central frequency. On the other hand, the LCLS-II average sub-pulse intensity is maximized at zero coherence width, followed by a slow but consistent decline until it reaches its asymptotic value. In the time domain of panel (b), we see that all parameter sets display similar qualitative behavior, with each average sub-pulse intensity curve monotonically increasing towards its asymptotic Fourier-limited value. Comparing the sub-pulse intensity distributions between the two panels (shaded areas), we see that broader distributions in the frequency domain are associated with narrower ones in the time domain, and vice versa, as a result of the uncertainty principle. We also see that the range of coherence widths required to reach the Fourier-limited asymptotic value is considerably larger in the time domain than in the frequency domain, as we observed in figure 2 for the FLASH case.
@@
Figure 3: (a-b) Simulated mean sub-pulse intensities (dotted) and sub-pulse modes (solid) as functions of the relative coherence width, defined as the ratio between the coherence width and average pulse bandwidth, for the LCLS-I (blue), LCLS-II (orange), and FLASH (purple) FEL parameter sets of table 1 in the (a) frequency and (b) time domains. For each data set, the upper and lower bounds of the shaded regions represent the 90th and 10th percentiles of the sub-pulse intensity distributions, respectively. In panel (a), we scaled the FLASH results for visual clarity. (c,d) Simulated average number of sub-pulses (curves) plotted against the relative coherence width for the LCLS-I (blue), LCLS-II (orange), and FLASH (purple) FEL parameter sets of table 1 in the (c) frequency and (d) time domains. For each curve, the shaded regions represent $\pm$1 standard deviation around mean value. (e-h) Histograms of the sub-pulse number distributions simulated for three exemplary values of the coherence width for the (e,f) LCLS-I (blue) and (g,h) FLASH (purple) parameter sets in the (e,g) frequency and (f,h) time domains.
@@
Next, we investigate the coherence width’s effects on the number of sub-pulses contained within a given VCM pulse. Figure 3 (c,d) compares how the average number of sub-pulses depends on the coherence width, simulated across all parameter sets and in both the (c) frequency and (d) time domains. In all cases, we see that the average number of sub-pulses is maximized at zero coherence width. Then, we find that increasing the coherence width results in the average number of sub-pulses monotonically decreasing towards one, just like we would expect from a fully-coherent pulse. In the FLASH and LCLS-I cases, we further find that the decrease follows an exponential law. Like for the sub-pulse intensities of figure 3 (a,b), we see that the distributions of sub-pulse numbers narrow with increasing coherence width – see the shaded areas at $\pm 1$ standard deviation around the average.
@@
Figure 3 (e-h) shows individual sub-pulse number distributions at three exemplary values of the coherence width simulated with the LCLS-I (top row) and FLASH (bottom row) parameter sets within both the frequency (left column) and time (right column) domains. At zero coherence width, we find that both parameter sets yield similar number distributions between the time and frequency domains, with a mean of $\approx 11$ sub-pulses for LCLS-I and $\approx 4$ sub-pulses for FLASH. As the coherence width increases well beyond the bandwidth of each parameter set, the distributions approach the single-sub-pulse limit for smaller coherence widths in the time domain than in frequency. For each panel (e-h), the leftmost distributions correspond to the minimal coherence width required such that at least 90% of VCM pulses contain a single sub-pulse, thereby reliably producing pulses that are effectively coherent. For the LCLS-I parameter set, this coherence width is 28.9 eV ($\approx 5.2$ times the average pulse bandwidth) in the frequency domain and 23.1 eV ($\approx 4.2$ times) in the time domain, whereas for FLASH it is 3.58 eV ($\approx 6.5$ times) in the frequency domain and 2.62 eV ($\approx 4.7$ times) in the time domain. Similar to the sub-pulse intensity, we attribute the differences in the minimum coherence widths required for convergence of the sub-pulse number in time vs frequency to the effect of increasing the coherence width on the shapes of the pulses in the time (narrow central peak with pedestals) and frequency (limited effect on the average pulse envelope) domains.
3.2 Joint Probability Distributions